# Supercomputing for High-speed Avoidance and Reactive Planning in Robots

### v1

@@ Abstract
This paper presents SHARP (Supercomputing for High-speed Avoidance and Reactive Planning), a proof-of-concept study demonstrating how high-performance computing (HPC) can enable millisecond-scale responsiveness in robotic control. While modern robots face increasing demands for reactivity in human–robot shared workspaces, onboard processors are constrained by size, power, and cost. Offloading to HPC offers massive parallelism for trajectory planning, but its feasibility for real-time robotics remains uncertain due to network latency and jitter. We evaluate SHARP in a stress-test scenario where a 7-DOF manipulator must dodge high-speed foam projectiles. Using a parallelized multi-goal A* search implemented with MPI on both local and remote HPC clusters, the system achieves mean planning latencies of 22.9 ms (local) and 30.0 ms (remote,  300 km away), with avoidance success rates of 84% and 88%, respectively. These results show that when round-trip latency remains within the tens-of-milliseconds regime, HPC-side computation is no longer the bottleneck, enabling avoidance well below human reaction times. The SHARP results motivate hybrid control architectures: low-level reflexes remain onboard for safety, while bursty, high-throughput planning tasks are offloaded to HPC for scalability. By reporting per-stage timing and success rates, this study provides a reproducible template for assessing real-time feasibility of HPC-driven robotics. Collectively, SHARP reframes HPC offloading as a viable pathway toward dependable, reactive robots in dynamic environments.
@@ I Introduction
Modern robots are increasingly expected to operate in unstructured and dynamic environments, often in close collaboration with humans. These scenarios demand not only accurate planning but also highly reactive control: a robot must perceive changes and adapt within tens of milliseconds to avoid unsafe or undesirable interactions. Traditionally, such responsiveness has relied on local computing platforms (commodity CPUs or GPUs embedded on the robot or on nearby workstations), which are constrained by size, power, and cost [1, 2]. As robots integrate more resource-intensive artificial intelligence (AI) and machine learning models, these platforms are reaching their limits, creating bottlenecks that threaten truly responsive human–robot interaction (HRI).
High-performance computing (HPC), or supercomputing, offers an alternative. HPC systems perform billions of computations per second and excel at large-scale graph search and optimization—problems central to robotics tasks such as inverse kinematics, trajectory planning, and collision avoidance [3, 4]. In principle, HPC can deliver solutions orders of magnitude faster than local compute, enabling behaviours that are more optimal, adaptive, and reactive.
The main obstacle is latency [5]. HPC resources are typically non-local and accessed via network connections that introduce delays. In time-critical robotics, where milliseconds matter, the central question is whether the raw compute advantage of HPC can overcome communication costs and still enable real-time action [6, 7].
@@ II-A High-Performance and Cloud Computing in Robotics
Robotics research has increasingly explored HPC to overcome the limitations of onboard processors for demanding tasks as robots often cannot carry high-end CPUs/GPUs needed for demanding algorithms [8]. ‘Cloud robotics’ paradigms offers an alternative by offloading computation to clusters, effectively giving robots access to scalable processing power [9]. Early frameworks like DAvinCi and Rapyuta exemplified this approach: DAvinCi integrated a Hadoop-based cluster with ROS to perform computationally intensive jobs (e.g. SLAM) in the cloud [10], while Rapyuta allowed robots to offload tasks and share a cloud-based knowledge repository (RoboEarth) to augment their local intelligence[11]. Since then, cloud-based HPC has supported core robotic tasks, often with significant performance gains [12].
However, off-site HPC introduces challenges. Network latency undermines real-time autonomy, and concerns remain about security and reliance on external infrastructure [13]. Recent frameworks like FogROS and FogROS2 address these concerns by allowing selective offloading: developers can designate which ROS nodes run locally versus in the cloud [8], [6]. In this architecture, a robot can offload heavy components (like SLAM, grasp planning, or motion planning) to cloud GPUs/CPUs. This hybrid approach has achieved order-of-magnitude speedups with modest overhead (0.5–1.2 s latency for 4–31× faster computation). This demonstrates that HPC resources (whether remote data centers or local clusters) can dramatically accelerate robotic algorithms, as long as the system is designed to tolerate the communication delays.
@@ II-B HPC and Cloud-based Planning for Robot Obstacle Avoidance
Trajectory planning and obstacle avoidance in complex environments remain computationally intensive, motivating HPC-based methods; traditional single-CPU planners often struggle to meet real-time requirements when the search space or number of obstacles grows large[14]. Parallel sampling-based approaches mitigate this: multiple RRT or RRT* instances can run concurrently, halting when the first finds a path [15]. Ichnowski et al. extend this to cloud serverless computing, launching many parallel planners as independent functions that internally exploit multi-core CPUs, enabling interactive planning for high-dimensional tasks [16]. This multi-level parallelism (threads within each function, and many functions in parallel) allowed motion plans to be found much faster than a single-core baseline, effectively approaching interactive planning even for high-dimensional manipulation tasks. The ability to scale out computation on demand – even if only for bursts – is a clear advantage of HPC in time-sensitive planning: a robot facing a difficult planning problem can temporarily enlist dozens of cores in the cloud to find a path in seconds rather than minutes.
@@ III Contributions
    We adapt a deterministic, voxelized A* search to HPC clusters via MPI, exploiting large-scale parallelism to guarantee bounded runtimes for evasive motion planning.
@@ IV-B1 Robot
We use a Franka Emika Panda, a 7-DOF manipulator. Its redundant configuration allows the robot to perform evasive manoeuvres while maintaining a desired position and orientation of the end-effector if allowed. The robot runs under ROS2 Humble with a real-time control frequency of 1 kHz (through the Franka Control Interface), ensuring fast and reliable command execution [17].
@@ IV-C System Workflow
On the HPC side, the voxel data is distributed across processes and evaluated in parallel. A multi-goal A* search identifies a feasible avoidance trajectory, which is then returned to the robot in compact waypoint form. The robot immediately executes this motion using a real-time controller, before resuming its monitoring state.
@@ V-B1 Graph Formulation
We store nodal information in four bit arrays: (1) node parents, (2) node distances from the starting node, (3) whether the node has been checked, and (4) whether it is invalid. As an example, discretizing the workspace into 1 cm3 voxels with a 50 cm robot radius and 8 elbow configurations requires only $\sim$1 MB of storage. Increasing the resolution to 1 mm3 voxels under the same conditions increases storage to $\sim$1 GB.
V-B2 Parallel Multi-Goal A*
We employ a parallelized multi-goal A* search adapted for MPI. Unlike RRT-based planners, which rely on random sampling and can require many iterations to converge [15], A* is deterministic and guarantees a feasible solution within bounded time—critical for real-time avoidance. Each process expands nodes from its local heap, distributes neighbours, and synchronizes with an MPI Allreduce to detect the first collision-free solution. The heuristic prioritizes minimal elbow displacement before end-effector distance, keeping evasive motions compact. Algorithm 1 summarizes the procedure. The heuristic guiding the search is defined as:
@@ V-C Trajectory Execution on Robot
Upon receiving waypoints, the robot interpolates them with the Ruckig library [18] to ensure smooth velocity and acceleration profiles. Execution proceeds under a 1 kHz proportional–derivative (PD) controller that converts tracking errors into motor torques. To reduce latency, the controller updates targets dynamically—advancing to the next waypoint once $\sim$20% of the current path is completed. After finishing an avoidance maneuver, the robot holds its final pose and resumes monitoring for subsequent projectiles.
@@ V-D Timing Breakdown
    Search: the core parallel A* traversal and trajectory identification.
@@ VI-A1 Blaster and Timing Budget
Average projectile speed at exit was $\sim 10~\text{m/s}$ with the blaster at $\sim 4.5~\text{m}$ from the robot, yielding a minimum time-of-flight of $\sim 450~\text{ms}$. For context, simple human visual reaction time is $\sim 200\text{–}250~\text{ms}$ [19]. This establishes a conservative envelope within which SHARP must complete sensing, offload, planning, return, and execution.
@@ VI-A3 Hardware
The robot manipulator was connected to a Dell G15 laptop (Ryzen 7 6800H) via Ethernet, allowing for use of the Franka Robot Interface for 1kHz communication [17]. For our experiments, we ran HPC computation on single compute node clusters both locally and remote (see Sec. IV-B4), and on the local, robot-connected laptop (Dell G15) for comparison. The HPC computation is able to run on more compute nodes; however, for initial testing, we use only one. Unfortunately, the laptop was unable to execute the full graph search at all due to a lack of sufficient memory and cores to run full tests.
@@ VI-B Local HPC Results
Table II presents the compute timings for local cluster runs. Local, round-trip cluster connections used throughout the trials were measured at less than 1 ms. On average, the end-to-end compute loop required 22.9 ms, well below human visual reaction thresholds ($\sim$250 ms) [19]. The majority of this time was spent in the search phase ($\sim$20 ms), with voxel unpacking averaging just over 2 ms and backtracing under 0.2 ms. Only one trial (L5) experienced a notable spike ($\sim$49.8 ms), but even this outlier remained substantially faster than human reflexes.
@@ VI-D Key Findings and Practical Envelope
We find that HPC planning fits comfortably within reaction budgets. Mean planning times were $22.9~\text{ms}$ (local) and $30.0~\text{ms}$ (remote), with rare spikes $55~\text{ms}$. Even accounting for short-haul RTT ($\sim 16~\text{ms}$), SHARP leaves substantial time for robot execution relative to $\sim 450~\text{ms}$ dart flight and $\sim 250~\text{ms}$ human reaction.
@@ VII-A Compute Perfomrnace and Latency in Context
Across local and short-haul remote trials, HPC planning completed in the tens of milliseconds, with search dominating and unpack/backtrace negligible (Tables II, IV). Projectile avoidance success on valid shots was similar for both clusters, averaging $84\%$ and $88\%$, respectively (Tables I, III). This indicates that, once an avoidance request is formed, HPC-side computation fits comfortably within typical reaction budgets, even when examined in context with other delays (e.g., sensing, message (de)serialization, cluster scheduling), leaving time for execution. The main exogenous factor is network RTT: sub-ms locally vs. $\sim$16 ms short-haul (and $\sim$60 ms intercontinental for reference). Our results suggest offloading remains viable while RTT stays in the tens-of-milliseconds regime, consistent with wired industrial Ethernet and emerging low-jitter wireless (5G/6G) [20], and well within a $\sim$250 ms human reaction envelope [19].
@@ VII-D Practical Deployments
At the scale of our demonstration ($\sim$17 MB), a single compute node was sufficient. The benefits of HPC become more pronounced as problem sizes grow: larger voxel grids, more elbow configurations, or fleets of collaborative robots quickly exceed onboard compute limits. Mobile platforms already encounter this during computationally intensive tasks such as SLAM, grasp planning, or high-dimensional motion planning. Compared to FogROS2, which reduced motion planning to 1.2 seconds on 96 cores, our millisecond-level results on 64 cores suggest that HPC could enable real-time SLAM and similar workloads [6].
In more complex scenarios, HPC parallelism scales naturally, allowing multi-robot systems to coordinate trajectories in dynamic environments without sacrificing responsiveness [9]. Established programming models such as MPI make it possible to expand planning and perception workloads without redesigning algorithms, and even the smallest system on the TOP500 list provides over 8,000 cores—far beyond the requirements of typical robotics workloads [21].
@@ Acknowledgement
We would like to thank Jacob Chrisholm for his support in the development of this system, as well as members of MItHRIL, CAESAR and Ingenuity Labs at Queen’s University for their roles in supporting this work.

### v3

@@ Abstract
This paper presents SHARP (Supercomputing for High-speed Avoidance and Reactive Planning), a proof-of-concept study demonstrating how high-performance computing (HPC) can enable millisecond-scale responsiveness in robotic control. While modern robots face increasing demands for reactivity in human–robot shared workspaces, onboard processors are constrained by size, power, and cost. Offloading to HPC offers massive parallelism for trajectory planning, but its feasibility for real-time robotics remains uncertain due to network latency and jitter. We evaluate SHARP in a stress-test scenario where a 7-DOF manipulator must dodge high-speed foam projectiles. Using a hash-distributed multi-goal A* search implemented with MPI on both local and remote HPC clusters, the system achieves mean planning latencies of 22.9 ms (local) and 30.0 ms (remote,  300 km away), with avoidance success rates of 84% and 88%, respectively. These results show that when round-trip latency remains within the tens-of-milliseconds regime, HPC-side computation is no longer the bottleneck, enabling avoidance well below human reaction times. The SHARP results motivate hybrid control architectures: low-level reflexes remain onboard for safety, while bursty, high-throughput planning tasks are offloaded to HPC for scalability. By reporting per-stage timing and success rates, this study provides a reproducible template for assessing real-time feasibility of HPC-driven robotics. Collectively, SHARP reframes HPC offloading as a viable pathway toward dependable, reactive robots in dynamic environments.
@@ I Introduction
Modern robots are increasingly expected to operate in unstructured and dynamic environments, often in close collaboration with humans. These scenarios demand not only accurate planning but also highly reactive control: a robot must perceive changes and adapt within tens of milliseconds to avoid unsafe or undesirable interactions. Traditionally, such responsiveness has relied on local computing platforms (commodity CPUs or GPUs embedded on the robot or on nearby workstations), which are constrained by size, power, and cost [AliOnRobots, Neuman2022TinyRobots]. As robots integrate more resource-intensive artificial intelligence (AI) and machine learning models, these platforms are reaching their limits, creating bottlenecks that threaten truly responsive human–robot interaction (HRI).
High-performance computing (HPC), or supercomputing, offers an alternative. HPC systems perform billions of computations per second and excel at large-scale graph search and optimization—problems central to robotics tasks such as inverse kinematics, trajectory planning, and collision avoidance [Rakita2019STAMPEDE:Kinematics, Natarajan2024PINSAT:Planning]. In principle, HPC can deliver solutions orders of magnitude faster than local compute, enabling behaviours that are more optimal, adaptive, and reactive.
The main obstacle is latency [Seisa2025Cloud-AssistedImplementation]. HPC resources are typically non-local and accessed via network connections that introduce delays. In time-critical robotics, where milliseconds matter, the central question is whether the raw compute advantage of HPC can overcome communication costs and still enable real-time action [IchnowskiFogROS2:2, Tahir2025EdgeSurvey].
@@ II-A High-Performance and Cloud Computing in Robotics
Robotics research has increasingly explored HPC to overcome the limitations of onboard processors for demanding tasks as robots often cannot carry high-end CPUs/GPUs needed for demanding algorithms [Chen2021FogROS:Deployment]. ‘Cloud robotics’ paradigms offers an alternative by offloading computation to clusters, effectively giving robots access to scalable processing power [Camargo-Forero2018TowardsComputing]. Early frameworks like DAvinCi and Rapyuta exemplified this approach: DAvinCi integrated a Hadoop-based cluster with ROS to perform computationally intensive jobs (e.g. SLAM) in the cloud [Arumugam2010DAvinCi:Robots], while Rapyuta allowed robots to offload tasks and share a cloud-based knowledge repository (RoboEarth) to augment their local intelligence[Mohanarajah2015Rapyuta:Platform]. Since then, cloud-based HPC has supported core robotic tasks, often with significant performance gains [Wan2016CloudIssues].
However, off-site HPC introduces challenges. Network latency undermines real-time autonomy, and concerns remain about security and reliance on external infrastructure [Du2017RobotComputing]. Recent frameworks like FogROS and FogROS2 address these concerns by allowing selective offloading: developers can designate which ROS nodes run locally versus in the cloud [Chen2021FogROS:Deployment], [IchnowskiFogROS2:2]. In this architecture, a robot can offload heavy components (like SLAM, grasp planning, or motion planning) to cloud GPUs/CPUs. This hybrid approach has achieved order-of-magnitude speedups with modest overhead (0.5–1.2 s latency for 4–31× faster computation). This demonstrates that HPC resources (whether remote data centers or local clusters) can dramatically accelerate robotic algorithms, as long as the system is designed to tolerate the communication delays.
@@ II-B HPC and Cloud-based Planning for Robot Obstacle Avoidance
Trajectory planning and obstacle avoidance in complex environments remain computationally intensive, motivating HPC-based methods; traditional single-CPU planners often struggle to meet real-time requirements when the search space or number of obstacles grows large[Abuelsamen2025IndustrialSystems]. Parallel sampling-based approaches mitigate this: multiple RRT or RRT* instances can run concurrently, halting when the first finds a path [Kuffner2000RRT-connect:Planning]. Ichnowski et al. extend this to cloud serverless computing, launching many parallel planners as independent functions that internally exploit multi-core CPUs, enabling interactive planning for high-dimensional tasks [Ichnowski2020FogComputing]. This multi-level parallelism (threads within each function, and many functions in parallel) allowed motion plans to be found much faster than a single-core baseline, effectively approaching interactive planning even for high-dimensional manipulation tasks. The ability to scale out computation on demand – even if only for bursts – is a clear advantage of HPC in time-sensitive planning: a robot facing a difficult planning problem can temporarily enlist dozens of cores in the cloud to find a path in seconds rather than minutes.
@@ III Contributions
    We adapt a deterministic, voxelized hash-distributed A* search to HPC clusters via MPI, exploiting large-scale parallelism to guarantee bounded runtimes for evasive motion planning.
@@ IV-B1 Robot
We use a Franka Emika Panda, a 7-DOF manipulator. Its redundant configuration allows the robot to perform evasive manoeuvres while maintaining a desired position and orientation of the end-effector if allowed. The robot runs under ROS2 Humble with a real-time control frequency of 1 kHz (through the Franka Control Interface), ensuring fast and reliable command execution [FrankaDocumentation].
@@ IV-C System Workflow
On the HPC side, the voxel data is distributed across processes and evaluated in parallel. A hash-distributed multi-goal A* search identifies a feasible avoidance trajectory, which is then returned to the robot in compact waypoint form. The robot immediately executes this motion using a real-time controller, before resuming its monitoring state.
@@ V-B1 Graph Formulation
We store nodal information in four bit arrays: (1) node parents, (2) node distances from the starting node, (3) whether the node has been checked, and (4) whether it is invalid. As an example, discretizing the workspace into 1 cm3 voxels with a 50 cm robot radius and 8 elbow configurations requires only $\sim$100 MB of storage. Increasing the resolution to 1 mm3 voxels under the same conditions increases storage to $\sim$100 GB.
V-B2 Hash-Distributed Multi-Goal A*
We employ a hash-distributed multi-goal A* search adapted for MPI. Unlike RRT-based planners, which rely on random sampling and can require many iterations to converge [Kuffner2000RRT-connect:Planning], A* is deterministic and guarantees a feasible solution within bounded time—critical for real-time avoidance. Each process expands nodes from its local heap, distributes neighbours, and synchronizes with an MPI Allreduce to detect the first collision-free solution. The heuristic prioritizes minimal elbow displacement before end-effector distance, keeping evasive motions compact. Algorithm 1 summarizes the procedure. The heuristic guiding the search is defined as:
@@ V-C Trajectory Execution on Robot
Upon receiving waypoints, the robot interpolates them with the Ruckig library [RuckigMachines] to ensure smooth velocity and acceleration profiles. Execution proceeds under a 1 kHz proportional–derivative (PD) controller that converts tracking errors into motor torques. To reduce latency, the controller updates targets dynamically—advancing to the next waypoint once $\sim$20% of the current path is completed. After finishing an avoidance maneuver, the robot holds its final pose and resumes monitoring for subsequent projectiles.
@@ V-D Timing Breakdown
    Search: the core hash-distributed A* traversal and trajectory identification.
@@ VI-A1 Blaster and Timing Budget
Average projectile speed at exit was $\sim 10~\text{m/s}$ with the blaster at $\sim 4.5~\text{m}$ from the robot, yielding a minimum time-of-flight of $\sim 450~\text{ms}$. For context, simple human visual reaction time is $\sim 180\text{–}200~\text{ms}$ [Reaction110800]. This establishes a conservative envelope within which SHARP must complete sensing, offload, planning, return, and execution.
@@ VI-A3 Hardware
The robot manipulator was connected to a Dell G15 laptop (Ryzen 7 6800H) via Ethernet, allowing for use of the Franka Control Interface for 1kHz communication [FrankaDocumentation]. For our experiments, we ran HPC computation on single compute node clusters both locally and remote (see Sec. IV-B4), and on the local, robot-connected laptop (Dell G15) for comparison. The HPC computation is able to run on more compute nodes; however, for initial testing, we use only one. Unfortunately, the laptop was unable to execute the full graph search at all due to a lack of sufficient memory and cores to run full tests.
@@ VI-B Local HPC Results
Table II presents the compute timings for local cluster runs. Local, round-trip cluster connections used throughout the trials were measured at less than 1 ms. On average, the end-to-end compute loop required 22.9 ms, well below human visual reaction thresholds ($\sim$200 ms) [Reaction110800]. The majority of this time was spent in the search phase ($\sim$20 ms), with voxel unpacking averaging just over 2 ms and backtracing under 0.2 ms. Only one trial (L5) experienced a notable spike ($\sim$49.8 ms), but even this outlier remained substantially faster than human reflexes.
@@ VI-D Key Findings and Practical Envelope
We find that HPC planning fits comfortably within reaction budgets. Mean planning times were $22.9~\text{ms}$ (local) and $30.0~\text{ms}$ (remote), with rare spikes $55~\text{ms}$. Even accounting for short-haul RTT ($\sim 16~\text{ms}$), SHARP leaves substantial time for robot execution relative to $\sim 450~\text{ms}$ dart flight and $\sim 200~\text{ms}$ human reaction.
@@ VII-A Compute Perfomrnace and Latency in Context
Across local and short-haul remote trials, HPC planning completed in the tens of milliseconds, with search dominating and unpack/backtrace negligible (Tables II, IV). Projectile avoidance success on valid shots was similar for both clusters, averaging $84\%$ and $88\%$, respectively (Tables I, III). This indicates that, once an avoidance request is formed, HPC-side computation fits comfortably within typical reaction budgets, even when examined in context with other delays (e.g., sensing, message (de)serialization, cluster scheduling), leaving time for execution. The main exogenous factor is network RTT: sub-ms locally vs. $\sim$16 ms short-haul (and $\sim$60 ms intercontinental for reference). Our results suggest offloading remains viable while RTT stays in the tens-of-milliseconds regime, consistent with wired industrial Ethernet and emerging low-jitter wireless (5G/6G) [Li2025ExploringNetwork], and well within a $\sim$200 ms human reaction envelope [Reaction110800].
@@ VII-D Practical Deployments
At the scale of our demonstration ($\sim$3 GB), a single compute node was sufficient. The benefits of HPC become more pronounced as problem sizes grow: larger voxel grids, more elbow configurations, or fleets of collaborative robots quickly exceed onboard compute limits. Mobile platforms already encounter this during computationally intensive tasks such as SLAM, grasp planning, or high-dimensional motion planning. Compared to FogROS2, which reduced motion planning to 1.2 seconds on 96 cores, our millisecond-level results on 64 cores suggest that HPC could enable real-time SLAM and similar workloads [IchnowskiFogROS2:2].
In more complex scenarios, HPC parallelism scales naturally, allowing multi-robot systems to coordinate trajectories in dynamic environments without sacrificing responsiveness [Camargo-Forero2018TowardsComputing]. Established programming models such as MPI make it possible to expand planning and perception workloads without redesigning algorithms, and even the smallest system on the TOP500 list provides over 8,000 cores—far beyond the requirements of typical robotics workloads [TOP500TOP500].
@@ Acknowledgement
We would like to thank members of MItHRIL, CAESAR and Ingenuity Labs at Queen’s University for their roles in supporting this work.